OpenStudy (anonymous):

find dy/dx expressed as a function of t for the given the parametric equations: x=cos^3(t) y=9sin^2(t) dy/dx=? not sure how to solve this problem, should i be plugging in values for t?

OpenStudy (freckles):

\[\frac{dy}{dx}=\frac{dy}{dt} \cdot \frac{dt}{dx} \text{ or } \frac{\frac{dy}{dt}}{\frac{dx}{dt}}\]

rvc (rvc):

differenitate x w.r.t.t and y w.r.t.t then

rvc (rvc):